OpenStudy (casodi):

Need help with working out the volume of a Tan(x) graph. My answer is always negative!

OpenStudy (casodi):

y=0 and x=pi/3 about the x-axis

OpenStudy (evoker):

So what are you integrating?

OpenStudy (casodi):

\[\pi*\int\limits_{0}^{\pi/3}(tanx)^2dx\]

OpenStudy (evoker):

ok that looks good

OpenStudy (casodi):

I was unsure whether the formula was correct

OpenStudy (evoker):

yeah since you are oscillating about the x-axis I would go with the integral of pi*y^2

OpenStudy (casodi):

after integrating tan^2x I get tanx-x. using that after substituting pi/3 I get a negative volume which I believe is impossible

OpenStudy (casodi):

not sure what im missing

OpenStudy (evoker):

hmm let me see, are you in radian mode on calc.

OpenStudy (casodi):

Dammit!!!

OpenStudy (casodi):

I completely forgot about that

OpenStudy (casodi):

I was in degrees

OpenStudy (evoker):

Yeah that would do it, I think you were good.

OpenStudy (casodi):

thanks for the help